OpenStudy (anonymous):

integrate of (x^4 + 1)/(x^2 + 1) dx is

OpenStudy (ghazi):

\[(x^4+1)(x^2+1)= x^6+x^4+x^2+1\] i have used law of distribution ..now you can integrate easily

OpenStudy (anonymous):

i not multiply, it's divide.

OpenStudy (anonymous):

\[\frac{x^4+1}{x^2+1}=\frac{x^4-1+2}{x^2+1}=\frac{(x^2-1)(x^2+1)+2}{x^2+1}\]make sense?

OpenStudy (anonymous):

so, how to integrate 2/(x^2 + 1) dx ?

OpenStudy (anonymous):

first long division. then use Partial fractions.

myininaya (myininaya):

How do you integrate 1/(x^2+1)?

myininaya (myininaya):

This is a memory thing. Or you could go the long way and get it .

OpenStudy (anonymous):

d/dx (arctan(x)=1/(1+x^2)

OpenStudy (anonymous):

i stil don't understand

OpenStudy (ghazi):

\[\frac{ d }{ dx }(\tan^{-1}x)= \frac{ 1 }{ 1+x^2 }\]

OpenStudy (ghazi):

\[\int\limits_{?}^{?} (\frac{ 1 }{ 1+x^2 })= \tan^{-1}x\]

OpenStudy (anonymous):

ok, thx for your help, i so appreciate :)

OpenStudy (ghazi):

:)